OpenStudy (anonymous):

John bowled 136 on his first game and 148 on his second game. Jerry had a total score of 483 for three games. What would John have to bowl on his third game to beat Jerry

OpenStudy (texaschic101):

136 + 148 = 284 (John's score so far) 483 - 284 = 199 (to tie Jerry's game. add 1 more to beat Jerry's game making John have to get 200 to beat Jerry. Understand ?